# Predicting and Understanding Turn-Taking Behavior in Open-Ended Group Activities in Virtual Reality

### v1

@@ 6. Limitations and Future Work
There are several limitations to our work. To start, we did not incorporate features related to the verbal transcript data. Future works should examine ways for extracting additional meaningful verbal features using transcripts, for example through extracting features with Linguistic Inquiry and Word Count (LIWC) (Tausczik and Pennebaker, 2010) for users at longer windows prior to the start of turn-taking behaviors. To balance model performance, complexity, and feature interpretation, we used summary statistics of motion data and chose standard machine learning models for predicting turn-taking behaviors. We imagine future work exploring other deep learning architectures such as graph neural networks (Wu et al., 2021b) and long short-term memory networks (Yu et al., 2019) to further capture the nuances of group dynamics. Since we mostly sampled features either immediately or at particular moments before turn-transition behaviors, another avenue for future work is in analyzing real-time synchronous performance of these models. Using turn-taking predictions, we imagine social VR platforms reducing miscommunication through filtering out irrelevant sounds (Akselrad et al., 2023) and moderate social interactions to mitigate verbal harassment (Freeman et al., 2022; Schulenberg et al., 2023).
In the dataset we examined, students gathered in preassigned groups of three or four and remained in the same groups across all four weeks using the same social VR platform. The avatars on this platform, as is the case in most social VR platforms, also lacked subtle characteristics such as facial expressions, which could alter how individuals interpret social cues when transitioning from face-to-face interactions to immersive ones. To fully evaluate the robustness of our predictive models, researchers should also examine scenarios using different social platforms with varying levels of immersion (Abdullah et al., 2021) and avatar representation (Smith and Neff, 2018), where the users’ group memberships could change across sessions (Salehi and Bernstein, 2018), and gather in groups with a greater variety of group sizes (Han et al., 2023). Future studies can also investigate differences in prediction performance using social VR platforms that allow for hand tracking in addition to hand controller tracking. A related limitation pertains to the dataset’s convenience sample of university students. Research should study demographics more representative of the general population to investigate the generalizability of our insights. Additionally, while VR facilitates interactions akin to face-to-face ones, medium-specific affordances such as the ability to participate in past social interactions (Wang et al., 2024b, 2023, 2021, 2020) can foster different social dynamics. In VR, there can be discrepancies in representations of virtual and physical motion. For example, in ENGAGE, a user can be physically seated while their virtual avatar stands in the virtual space. Future research should therefore examine how these factors contribute to different turn-taking behavior patterns when compared to other digital platforms and face-to-face interactions.
Finally, we focused on prediction tasks related to three of four categories of turn-transition behaviors, excluding the category of backchanneling. As backchanneling and overlap speech make up a key component to social dynamics (Schegloff, 2000) and can have unique consequences in CMC given their latencies and the diminished representation of nonverbal behaviors (Smith and Neff, 2018; Sheng, 2021; Seuren et al., 2021), works can also investigate overlapping speech, for example creating approaches and mechanisms for predicting and reducing unwanted interruptions.
@@ 7. Conclusions
Being able to predict turn-taking behaviors in VR affords opportunities for understanding how social interactions unfold in immersive settings and enables systems to administer personalized and adaptive intervention. In this work, we studied the turn-taking behaviors of student groups engaging in open-ended group activities in VR over a period of four weeks. We predicted turn-taking behaviors using features describing individual and group characteristics as well as those extracted from VR tracking data concerning speech-related behaviors, egocentric motion, and dyadic and group relationships. We found that gradient boosting classifiers achieved the best performance, considerably better than prediction by chance. Analyses of these models revealed that listener personality, group size, group gaze, and the listener and previous speaker’s head pitch, head y-axis position, and left hand y-axis position were key features affecting prediction and performance. Our results suggested that these features are reliable indicators in novel social settings, as the models were robust when evaluated on unseen activities, weeks, and groups. Taken together, our work contributed a better understanding of how tracking data and individual and group characteristics can predict turn-taking behaviors in social VR settings. We believe that our insights will motivate further research on modeling social dynamics, and support practitioners to use behavioral predictions to deliver real-time adaptive assistance in facilitating effective immersive social interactions.
@@ Acknowledgements.
We thank Cindy Chen, Mark Roman Miller, Monique Santoso, and Yujie Tao for their feedback. This work was supported by the Sozo Graduate Fellowship.References
@@
- Wang et al. (2020) Cheng Yao Wang, Mose Sakashita, Upol Ehsan, Jingjin Li, and Andrea Stevenson Won. 2020. Again, Together: Socially Reliving Virtual Reality Experiences When Separated. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3313831.3376642
- Wang et al. (2021) Cheng Yao Wang, Sandhya Sriram, and Andrea Stevenson Won. 2021. Shared Realities: Avatar Identification and Privacy Concerns in Reconstructed Experiences. Proc. ACM Hum.-Comput. Interact. 5, CSCW2, Article 337 (oct 2021), 25 pages. https://doi.org/10.1145/3476078
@@ Appendix A Descriptions of Weekly Group Activities
4 | Education | • Consider a target audience/population (e.g., students of a certain age group, students with a certain learning disability, older students). • Consider a goal (e.g., retaining factual information, having students experience something) • Consider a topic of interest (e.g., language, STEM, social skills) • Collaboratively work with your group members to create a meditation room or a safe space using any of the ENGAGE tools (e.g., 3D pen, IFX, sticky notes) |
@@ Appendix B Feature Importance for All Features
![Refer to caption](extracted/5634250/figures/feature-importance-full-v2.png)

### v2

@@ 6. Limitations and Future Work
Our work has several limitations. To start, we did not interpret features related to prior speech sequences or incorporated features related to verbal transcripts. Works should examine how virtual speech sequences differ from face-to-face ones and investigate how they predict turn-taking dynamics. Researchers should also extract additional verbal features using transcripts, for example features using Linguistic Inquiry and Word Count (Tausczik and Pennebaker, 2010). Relatedly, though we did not encode virtual context characteristics, works should sample turn-taking behaviors in drastically-different virtual spaces and curate approaches for parameterizing them. Researchers should also explore whether incorporating gaze-related features enhances model performance since our dataset did not contain them. To balance model performance, complexity, and interpretability, we used summary statistics and standard machine learning models. We imagine works exploring other deep learning architectures (Wu et al., 2021b; Yu et al., 2019) to capture additional nuances of group dynamics. Another avenue of future work lies in benchmarking the real-time performance of these models and leveraging predictive models to reduce miscommunication (Akselrad et al., 2023) and mitigate verbal harassment (Freeman et al., 2022; Schulenberg et al., 2023).
In our dataset, students gathered in fixed groups using the same social VR platform and embodied avatars that lacked nonverbal cues such as facial expressions. To fully evaluate model robustness, researchers should examine different social platforms with varying levels of avatar representation (Smith and Neff, 2018) and immersion (Abdullah et al., 2021), change group membership over time (Salehi and Bernstein, 2018), and vary group sizes (Han et al., 2023). As noted in Section 5.1.1, differences in group sizes beyond the scope of our paper could yield drastic differences in social dynamics and lower the applicability of our approach. Different activity types, for example those varying in virtual mobility (e.g., scavenger hunt vs. discussion), could elicit similar concerns. Future research should therefore probe the limits to improving performance through naively scaling up data and sampling across a more diverse set of group interactions (e.g., group size, activity type). It may also be necessary to consider different machine learning architectures for modeling certain group activities, for example training expert sub-models to handle different social scenarios before integrating them into for making predictions (Lin et al., 2024). Doing so will not only suggest technical improvements to the applicability of predicting turn-taking behaviors but also built toward a better understanding of how VR contexts influence social dynamics.
One other limitation pertains to the dataset’s convenience sample of university students. To investigate the generalizability of our insights, researcher should study more representative demographics. Additionally, discrepancies between virtual and physical motion (e.g., physically seated users with standing virtual avatars) warrant scrutiny for their impact on turn-taking behaviors. Finally, while we examined three turn-transition categories, we excluded backchanneling. As backchanneling and overlap speech make up a key component to social dynamics (Schegloff, 2000) and have unique challenges in CMC due to latencies and diminished nonverbal behaviors (Smith and Neff, 2018; Sheng, 2021; Seuren et al., 2021), works should investigate overlapping speech, for example through predicting unwanted interruptions.
@@ 7. Conclusions
Being able to predict turn taking in VR affords opportunities for understanding immersive social interactions and enables systems to administer support and intervention. In this work, we studied turn-taking behaviors of student engaging in open-ended group activities in VR over four weeks. We predicted turn-taking behaviors using features describing individual and group characteristics and extracted features concerning speech-related behaviors, egocentric motion, and dyadic and group relationships. We found that gradient boosting classifiers achieved the best performance, considerably better than prediction by chance. Additional analysis revealed that listener personality, group size, group visual attention, and listener and previous speaker’s head pitch, head y-axis position, and left hand y-axis position were key features affecting performance and predictions. Our results suggest that these features are reliable indicators, as models were robust when evaluated on unseen activities, weeks, and groups. Taken together, our work contributed a better understanding of how tracking data and individual and group characteristics can predict VR turn taking. We believe our insights will motivate research on modeling social dynamics, and support practitioners to use behavioral predictions to deliver assistance and training in facilitating immersive social interactions.
@@ Acknowledgements.
Acknowledgements withheld for the review process.References
@@
- Bradley (1997) Andrew P Bradley. 1997. The use of the area under the ROC curve in the evaluation of machine learning algorithms. Pattern recognition 30, 7 (1997), 1145–1159.
- Brough et al. (2007) John E Brough, Maxim Schwartz, Satyandra K Gupta, Davinder K Anand, Robert Kavetsky, and Ralph Pettersen. 2007. Towards the development of a virtual environment-based training system for mechanical assembly operations. Virtual reality 11 (2007), 189–206.
@@
- Couronné et al. (2018) Raphael Couronné, Philipp Probst, and Anne-Laure Boulesteix. 2018. Random forest versus logistic regression: a large-scale benchmark experiment. BMC bioinformatics 19 (2018), 1–14.
- Degutyte and Astell (2021) Ziedune Degutyte and Arlene Astell. 2021. The role of eye gaze in regulating turn taking in conversations: a systematized review of methods and findings. Frontiers in Psychology 12 (2021), 616471.
@@
- Dormann et al. (2013) Carsten F Dormann, Jane Elith, Sven Bacher, Carsten Buchmann, Gudrun Carl, Gabriel Carré, Jaime R García Marquéz, Bernd Gruber, Bruno Lafourcade, Pedro J Leitão, et al. 2013. Collinearity: a review of methods to deal with it and a simulation study evaluating their performance. Ecography 36, 1 (2013), 27–46.
@@
- Freeman et al. (2002) Nancy L Freeman, Adrienne Perry, and James M Bebko. 2002. Behaviour is communication: Nonverbal communicative behaviour in students with autism and instructors’ responsivity. Journal on Developmental Disabilities 9, 2 (2002), 145–155.
@@
- Greenwell et al. (2018) Brandon M Greenwell, Bradley C Boehmke, and Andrew J McCarthy. 2018. A simple and effective model-based variable importance measure. arXiv preprint arXiv:1805.04755 (2018).
@@
- Hall et al. (1999) Cathy W Hall, Andrea D Peterson, Raymond E Webster, Larry M Bolen, and Michael B Brown. 1999. Perception of nonverbal social cues by regular education, ADHD, and ADHD/LD students. Psychology in the Schools 36, 6 (1999), 505–514.
@@
- Hare (1981) A Paul Hare. 1981. Group size. American Behavioral Scientist 24, 5 (1981), 695–708.
@@
- Huang et al. (2021) Gaoping Huang, Xun Qian, Tianyi Wang, Fagun Patel, Maitreya Sreeram, Yuanzhi Cao, Karthik Ramani, and Alexander J. Quinn. 2021. AdapTutAR: An Adaptive Tutoring System for Machine Tasks in Augmented Reality. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 417, 15 pages. https://doi.org/10.1145/3411764.3445283
@@
- Jun et al. (2022) Hanseul Jun, Mark Roman Miller, Fernanda Herrera, Byron Reeves, and Jeremy N. Bailenson. 2022. Stimulus Sampling With 360-Videos: Examining Head Movements, Arousal, Presence, Simulator Sickness, and Preference on a Large Sample of Participants and Videos. IEEE Transactions on Affective Computing 13, 3 (2022), 1416–1425. https://doi.org/10.1109/TAFFC.2020.3004617
@@
- Koiso et al. (1998) Hanae Koiso, Yasuo Horiuchi, Syun Tutiya, Akira Ichikawa, and Yasuharu Den. 1998. An analysis of turn-taking and backchannels based on prosodic and syntactic features in Japanese map task dialogs. Language and speech 41, 3-4 (1998), 295–321.
@@
- Lin et al. (2024) Bin Lin, Zhenyu Tang, Yang Ye, Jiaxi Cui, Bin Zhu, Peng Jin, Junwu Zhang, Munan Ning, and Li Yuan. 2024. Moe-llava: Mixture of experts for large vision-language models. arXiv preprint arXiv:2401.15947 (2024).
@@
- Lundberg (2017) Scott Lundberg. 2017. A unified approach to interpreting model predictions. arXiv preprint arXiv:1705.07874 (2017).
- Lundberg et al. (2018) Scott M Lundberg, Gabriel G Erion, and Su-In Lee. 2018. Consistent individualized feature attribution for tree ensembles. arXiv preprint arXiv:1802.03888 (2018).
@@
- Sato et al. (2023) Arissa J. Sato, Zefan Sramek, and Koji Yatani. 2023. Groupnamics: Designing an Interface for Overviewing and Managing Parallel Group Discussions in an Online Classroom. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 701, 18 pages. https://doi.org/10.1145/3544548.3581322
@@
- Sullivan and Feinn (2012) Gail M Sullivan and Richard Feinn. 2012. Using effect size—or why the P value is not enough. Journal of graduate medical education 4, 3 (2012), 279–282.
@@
- Wells and Windschitl (1999) Gary L Wells and Paul D Windschitl. 1999. Stimulus sampling and social psychological experimentation. Personality and Social Psychology Bulletin 25, 9 (1999), 1115–1125.
@@
- Yang and Berdine (2017) Shengping Yang and Gilbert Berdine. 2017. The receiver operating characteristic (ROC) curve. The Southwest Respiratory and Critical Care Chronicles 5, 19 (2017), 34–36.
@@ Appendix A Descriptions of Weekly Group Activities
4 | Education | • Consider a target audience/population (e.g., students of a certain age group, students with a certain learning disability, older students). • Consider a goal (e.g., retaining factual information, having students experience something) • Consider a topic of interest (e.g., language, STEM, social skills) • Empathize, define, ideate, and prototype an application tailored to your audience, goal, and topic. Have a member of your group test out/role-play a student using the application |
@@ Appendix B Feature Importance for All Features
Appendix C Wald Tests for Logistic Regression Models
We present results from joint Wald-tests ran on the trained logistic regression models, specifically the p values from the joint hypothesis tests that all features related to a specific construct do not significantly contribute to the predictive model. The orders of the feature groups and constructs follow that of Appendix B. Features with statistical significance, evaluated at , are bolded.
[b]
Turn Taking vs. | Next Speaker | Timing of | |
Continuing Speech | Prediction | Turn Taking | |
User has spoken | ¡.001 | ¡.001 | .264 |
User turn count | .046 | ¡.001 | .004 |
Prior speech sequence | ¡.001 | ¡.001 | ¡.001 |
User prev. turn time | ¡.001 | .046 | ¡.001 |
[b]
Turn Taking vs. | Next Speaker | Timing of | |
Continuing Speech | Prediction | Turn Taking | |
Ref. personality | .020 | ¡.001 | ¡.001 |
Group personality | .189 | ¡.001 | .191 |
Group size | .001 | ¡.001 | .646 |
User personality | .001 | ¡.001 | .189 |
[b]
Turn Taking vs. | Next Speaker | Timing of | |
Continuing Speech | Prediction | Turn Taking | |
Dyad IPD | .004 | .066 | .095 |
Dyad shared space | .177 | .096 | .628 |
Dyad ref. to user attn. | .846 | ¡.001 | .057 |
Dyad user to ref. attn. | .246 | ¡.001 | .087 |
[b]
Turn Taking vs. | Next Speaker | Timing of | |
Continuing Speech | Prediction | Turn Taking | |
Group user IPD | ¡.001 | .588 | .276 |
Ref. to group attn. | ¡.001 | ¡.001 | .016 |
User to group attn. | ¡.001 | .298 | .002 |
Group ref. IPD | .009 | .347 | .056 |
Group user shared space | .017 | .982 | .332 |
Group ref. shared space | ¡.001 | .784 | .292 |
Group to user attn. | ¡.001 | .080 | .007 |
Group to ref. attn. | ¡.001 | ¡.001 | .070 |
[b]
Turn Taking vs. | Next Speaker | Timing of | |
Continuing Speech | Prediction | Turn Taking | |
Ref. rh rot pitch | .028 | .964 | .297 |
Ref. lh rot pitch | .195 | .993 | .014 |
Ref. rh pos x | .143 | .997 | .007 |
User lh pos x | .015 | .181 | .566 |
Ref. head pos z | .018 | .588 | .505 |
Ref. rh rot yaw | ¡.001 | .984 | .187 |
Ref. head pos x | .154 | .541 | .190 |
Ref. head rot yaw | .202 | .823 | .277 |
Ref. lh rot roll | .252 | .991 | .766 |
Ref. lh rot yaw | .470 | .967 | .197 |
Ref. rh pos y | .493 | .992 | .103 |
User rh rot yaw | .002 | .017 | .173 |
User head rot yaw | .001 | .590 | .343 |
Ref. head rot roll | .058 | .947 | .191 |
Ref. lh pos x | .007 | .994 | .027 |
User lh rot roll | .961 | .316 | .043 |
User rh rot pitch | .564 | .225 | .665 |
User lh rot yaw | .012 | .052 | .035 |
Ref. rh pos z | .646 | .999 | .149 |
User rh pos z | .080 | .816 | .595 |
User lh rot pitch | .110 | .161 | .440 |
Ref. lh pos z | .198 | .977 | .057 |
User lh pos z | .185 | .246 | .021 |
User head pos x | .111 | .086 | .054 |
Ref. rh rot roll | ¡.001 | .971 | .090 |
User rh pos x | .048 | .215 | .558 |
User head pos z | .719 | .052 | .219 |
User head rot roll | .003 | ¡.001 | ¡.001 |
User rh rot roll | .002 | ¡.001 | .027 |
User rh pos y | .035 | .082 | .230 |
Ref. head pos y | ¡.001 | .605 | ¡.001 |
Ref. lh pos y | .102 | .942 | .213 |
User lh pos y | .002 | .027 | .087 |
Ref. head rot pitch | ¡.001 | .640 | .040 |
User head pos y | ¡.001 | .015 | ¡.001 |
User head rot pitch | ¡.001 | ¡.001 | ¡.001 |
Appendix D Additional Results on Model Predictions
In this section, we present additional analyses on prediction performance across sessions, groups, and weeks. To do this, we first randomly split each of the three datasets, one for each task, into a training set with 90% of the data and a testing set containing the remaining 10%. Then, we built models on the training data and evaluated them on the testing set in three ways: across sessions, groups, and weeks. Specifically, when evaluating across weeks, we calculated the average AUC and standard error given the trained models’ performances across four test sets, each containing the testing data from a specific week. We repeated this procedure for extracting the AUC and standard errors of the models across sessions and groups. Note that the reported average AUC here gives equal weight to each session, group, and week, which differs from the cross-validation evaluations presented in Section 4.1.
[b]
Performance Metrics | |||
Prediction Model | By Session | By Group | By Week |
Logistic Regression | 0.70 (0.02) | 0.71 (0.02) | 0.73 (0.01) |
MLP Classifier | 0.71 (0.02) | 0.70 (0.02) | 0.72 (0.02) |
Random Forest Classifier | 0.76 (0.01) | 0.77 (0.01) | 0.79 (0.01) |
Gradient Boosting Classifier | 0.80 (0.01) | 0.81 (0.01) | 0.82 (0.01) |
[b]
Performance Metrics | |||
Prediction Model | By Session | By Group | By Week |
Logistic Regression | 0.72 (0.02) | 0.72 (0.02) | 0.72 (0.01) |
MLP Classifier | 0.72 (0.02) | 0.72 (0.02) | 0.72 (0.01) |
Random Forest Classifier | 0.75 (0.02) | 0.75 (0.02) | 0.75 (0.00) |
Gradient Boosting Classifier | 0.75 (0.02) | 0.76 (0.02) | 0.77 (0.00) |
[b]
Performance Metrics | |||
Prediction Model | By Session | By Group | By Week |
Logistic Regression | 0.61 (0.02) | 0.62 (0.02) | 0.62 (0.01) |
MLP Classifier | 0.64 (0.01) | 0.64 (0.02) | 0.64 (0.021 |
Random Forest Classifier | 0.69 (0.01) | 0.68 (0.02) | 0.69 (0.01) |
Gradient Boosting Classifier | 0.72 (0.02) | 0.71 (0.02) | 0.73 (0.01) |
Appendix E Additional Results on Predicting Continuing Speech and Next Speaker
While our formulation of turn-taking behavior prediction tasks regarded turn-taking behavior vs. continuing speech and next speaker prediction as two individual tasks, we further present an alternative task that aims to predict speaking intentions more generally. This task holds practical importance as it addresses the broader question of forecasting who will speak next, be it speaker continuation or a switch to a new speaker. Importantly, this task differs from that of turn-taking behavior vs. continuing as the former seeks to distinguish between upcoming speakers and users who will not speak, whereas the latter focuses on distinguishing between two types of upcoming speakers.
Specifically, we defined positive samples as those sampled immediately preceding a continuing speech or clean turn taking speech event. For clean turn-taking behaviors, we chose the main user as the upcoming speaker and the previous speaker as the reference user. For samples corresponding to continuing speech, we chose the upcoming speaker as both the main and reference user. Negative samples were selected also at moments immediately preceding either a clean turn taking or continuing speech event. For the negative samples, we set the reference user as the previous speaker and the main user as one of the users who will not be the upcoming speaker. Across all samples, when the main user and the reference user are chosen to be the same user, we set the dyad-related features to zero after standardization. We down-ampled the training data such that there is an equal number of positive and negative data. Similar to Section 4.1, we report in Table 15 the AUC for the four machine learning models across the performance metrics introduced in Section 3.5.
We extended our binary classification models to multiclass prediction. Specifically, we evaluated each instance before a turn-taking or continuing speech event to predict which group member, including the current speaker, would speak next. Unlike our previous individual-based predictions, this formulation focuses on group-level predictions. Using the trained binary models from the previous step, we evaluated each user in the scene (i.e., chosen user as the main user, previous speaker as the reference user) and selected the user with the highest probability of a positive label as the predicted next speaker. Comparing the predictions with the true next speakers, we calculated the testing data accuracy and further reported separately the accuracies of the testing data with three-person groups and four-person groups. When presenting accuracies for cross-validation evaluations, we also report their standard errors across all folds. Tables 16–18 summarize our results.
[b]
Performance Metrics | ||||
Prediction Model | Sessioncv | Groupcv | Weekcv | Week 4 |
Logistic Regression | 0.74 (0.01) | 0.73 (0.01) | 0.74 (0.01) | 0.73 |
MLP Classifier | 0.75 (0.00) | 0.74 (0.01) | 0.74 (0.01) | 0.74 |
Random Forest Classifier | 0.79 (0.00) | 0.79 (0.01) | 0.79 (0.01) | 0.79 |
Gradient Boosting Classifier | 0.82 (0.00) | 0.81 (0.01) | 0.82 (0.00) | 0.81 |
[b]
Performance Metric: Sessioncv | |||
Prediction Model | All | Three-Person | Four-Person |
Logistic Regression | 0.51 (0.01) | 0.54 (0.01) | 0.48 (0.01) |
MLP Classifier | 0.54 (0.01) | 0.57 (0.01) | 0.51 (0.01) |
Random Forest Classifier | 0.57 (0.01) | 0.60 (0.01) | 0.54 (0.01) |
Gradient Boosting Classifier | 0.62 (0.01) | 0.65 (0.01) | 0.59 (0.01) |
[b]
Performance Metric: Groupcv | |||
Prediction Model | All | Three-Person | Four-Person |
Logistic Regression | 0.51 (0.01) | 0.53 (0.02) | 0.49 (0.02) |
MLP Classifier | 0.52 (0.01) | 0.55 (0.01) | 0.50 (0.02) |
Random Forest Classifier | 0.57 (0.01) | 0.58 (0.02) | 0.55 (0.02) |
Gradient Boosting Classifier | 0.61 (0.01) | 0.63 (0.01) | 0.59 (0.01) |
[b]
Performance Metric: Weekcv | |||
Prediction Model | All | Three-Person | Four-Person |
Logistic Regression | 0.51 (0.01) | 0.54 (0.01) | 0.48 (0.01) |
MLP Classifier | 0.54 (0.01) | 0.57 (0.01) | 0.49 (0.00) |
Random Forest Classifier | 0.57 (0.01) | 0.60 (0.01) | 0.54 (0.00) |
Gradient Boosting Classifier | 0.62 (0.01) | 0.65 (0.01) | 0.58 (0.00) |
As shown in Table 15, gradient boosting classifiers yielded the highest performance with an AUC of 0.81–0.82 across the four performance metrics. Similar to results presented in Section 4.1, the random forest classifiers yielded the second best performance with an AUC of 0.79, followed by the MLP classifiers and the logistic regressions at 0.74–0.75 AUC and 0.73–0.74 AUC. Tables 16–18 revealed similar trends, namely that the gradient boasting classifiers yielded the highest accuracy across all evaluation metrics (i.e., 59–65%). We also found that the accuracies among three-person groups are higher than that for four-person groups. For reference, random chance predictions would result in an accuracy of 33.33% for three-person groups and 25.00% for four-person groups.